UT 20130904 Technical Observing - AGw Calibration

Observer: MEdwards (Tucson) JHill (LBTO)
Telescope Operator: DGonzalez (LBTO)
Software Support: DCox (Tucson)
AO Support: DMiller (LBTO)
Telescope Support: JLittle, KNewton (LBTO)
Instrument Support: JPower, JMorris (LBTO)

Summary

Closed at the very beginning of the night for clouds. Tried to open and start transform in 0.8" seeing but closed quickly again for clouds

Details

01:45 David is leery of cloud cover esp. w/rain predicted in the forecast. Staying closed for now. LUCI software opened and ready to go; dark looks fine and transfers correctly to /newdata

02:06 Clouds have cleared and DGH has decided it is safe to open. First DIMM seeing is actually good ~ 0.8". We'll start with LUCI transformation which requires 0.8" or better. Choosing Stone_L1: 18:58:45.907 +00:03:14.96

02:30 LUCI transformation data-taking underway. Thanks to Doug M. for the advice and help getting started!

02:50 Happily taking transform data -- so far so good.

03:00 Unhappily closing for clouds

05:42 Opening again. Waiting to see GCS. Looks like it's 0.8" so still pretty good.

06:00 Finding new Stone field since old one is setting. Choosing Stone_O -- 21:47:54.022 +00:00:01.79. Starting at X = 1.187. Doing preset first for collimation before starting transform_collect

06:05 Starting transform_collect. Star on LUCI image looks like 120,000 counts with dit = 4 and ndit = 3. GCS reads ~0.6.

06:48 Clouds again. We were unfortunately about halfway through so that is a bummer. Seeing degraded in last few images, but was very good 0.6-0.8" until that point. Waiting for a while to decide what to do.

07:12 Thought about opening for a minute but decided against it. Still waiting it out.

07:50 Probably unlikely we are going to get 1.5-2 solid hours with no clouds now with LUCI for the transformation. Switching to LBC in hopes of getting something in -- if we open at all.

08:00 Opened with LBC ready. DGH driving from LBTO; MLE trains from Tucson

08:30 DGH successfully runs dofpia and lbcrangbal on WT10_240. Seeing is about 1"

08:50 DGH slews to NGC101A1. Runs dofpia again on target

09:00 DGH takes a few images of NGC101A1

09:15 Michelle turns it over for DGH to run superfoc sequences for tilt and filter offsets!

09:30 David called Kellee because it appeared the CMU had crashed. Kellee logged in to the console and watched when David power-cycled. It came up ok, different from when it crashed a few days ago. There was nothing in the system log to explain why it crashed.

-- MichelleEdwards - 07 Sep 2013
Topic revision: r5 - 09 Sep 2013, KelleeSummers
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback